please check my equation and let me know if I'm missing a step. Thank you
3(6-4x)=2(-6x+9)
18-4x =-4x+18
+4 +4
18 = 18

Answer by Fombitz(13828)

3(6-4x)=2(-6x+9)
18-4x =-4x+18<--- 3*4x=12x not 4x on the left hand side, 2(-6x)=-12x not -4x on the right hand side
+4 +4
18 = 18